# A non-degeneracy theorem for interacting electrons in one dimension

### v1

@@ 6.1. Unique continuation along the boundary
Let $\Gamma\subset\partial I_{N}$, and let $\Psi$ be the unique ground-state of the self-adjoint realization of $a_{v,w}$ with form domain $\mathrm{H}^{1}(I_{N};\Gamma)\cap\mathcal{H}_{N}$. Then $\gamma\Psi$ can not vanish identically on a relatively open set
@@
Moreover, if $\alpha\in\mathbb{R}\setminus\{0\}$ satisfies (2.4), then the same holds for the ground-state of the self-adjoint realization with form domain $\mathrm{H}^{1}_{\alpha}(I_{N};\Gamma)\cap\mathcal{H}_{N}$.
Proof of Theorem 6.1 in the case of local BCs.
@@
is continuous. We can thus define the extended self-adjoint operator $H_{\Omega}$ via the sesquilinear form
$\widetilde{a}_{v,w,\Omega}(\Psi,\Phi)$ $\coloneqq\int_{\Omega\setminus S_{N}}\overline{\nabla\Psi(x)}\bm{\cdot}\nabla\Phi(x)\mathrm{d}x+\widetilde{a}_{v,w}(R\Psi,R\Phi)$
$=\int_{\Omega}\overline{\nabla\Psi(x)}\bm{\cdot}\nabla\Phi(x)\mathrm{d}x+(T^{\#}v)(R\Psi,R\Phi)+(T^{\#}w)(R\Psi,R\Phi),\quad\Psi,\Phi\in\mathrm{H}^{1}(\Omega;\Gamma_{S}).$ (6.1)
@@
$\widetilde{a}_{v,w,\Omega}(\widetilde{\Psi},\Phi)=\int_{\Omega}\overline{\nabla\widetilde{\Psi}}\bm{\cdot}\nabla\Phi\mathrm{d}x+(T^{\#}(v+w))(R\widetilde{\Psi},R\Phi)=\widetilde{a}_{v,w}(\widetilde{\Psi},R\Phi)=\lambda\langle\widetilde{\Psi},R\Phi\rangle=\lambda\langle\widetilde{\Psi},\Phi\rangle,$
where $\lambda$ is the ground-state energy of $\widetilde{H}_{N}(v,w;\Gamma)$. In particular, $\widetilde{\Psi}$ is an eigenfunction of $H_{\Omega}$.
Next, notice that the pullback $R^{\#}V$ of any bounded multiplicative function $V\in\mathrm{L}^{\infty}(S_{N})$ is simply the extension of $V$ by zero to $\Omega$, so in particular, a bounded function in $\mathrm{L}^{\infty}(\Omega)$. Consequently, the pullback via $R$ of any sequence of bounded functions in $S_{N}$ approximating the quadratic form $T^{\#}(v+w)$ in $\mathrm{H}^{1}$ (as in (5.10)) yields a sequence of bounded functions approximating $R^{\#}T^{\#}(v+w)$. In particular, the operator $H_{\Omega}$ defined via (6.1) satisfies the assumptions of Theorem 5.3, and therefore, has a unique and a.e. strictly positive ground-state $\Psi_{\Omega}$.
This now yields a contradiction for the following reason. As we have shown that $\widetilde{\Psi}$ is an eigenfunction of $H_{\Omega}$, it must be either orthogonal or proportional to $\Psi_{\Omega}$. However, $\widetilde{\Psi}$ cannot be orthogonal to $\Psi_{\Omega}$ as they are both non-negative (up to a global phase), and $\widetilde{\Psi}$ cannot be parallel to $\Psi_{\Omega}$ as the former vanishes on $\Omega\setminus S_{N}$ while the later is (a.e.) strictly positive in $\Omega$. This completes the proof. ∎
@@
Let $\alpha\in\mathbb{R}\setminus\{0\}$ and $\Psi\in\mathrm{H}^{1}_{\alpha}(I_{N};\Gamma)\cap\mathcal{H}_{N}$ be an eigenfunction of $H_{N}^{\alpha}(v,w;\Gamma)$ with eigenvalue $\lambda$. Then for any $f\in\mathrm{H}^{1/2}(\partial I_{N})$, we define the Neumann trace $\partial_{\nu}\Psi$ as
$(\partial_{\nu}\Psi)(f)=\int_{I_{N}}\overline{\nabla\Psi(x)}\bm{\cdot}\nabla F(x)+v(\rho_{\Psi,P_{N}F})+w(\rho_{2,\Psi,P_{N}F})-\lambda\langle\Psi,F\rangle$ (6.2)
@@
$(P_{N}F)(x)=\frac{1}{N!}\sum_{\sigma\in\mathcal{P}_{N}}\mathrm{sgn}(\sigma)F(\sigma x).$
It is not hard to see that $\partial_{\nu}\Psi(f)$ is independent of the extension of $f$. Indeed, if we have two $\mathrm{H}^{1}$-extensions $F$ and $F^{\prime}$, then $F-F^{\prime}\in\mathrm{H}^{1}_{0}(I_{N})$, and therefore $P_{N}(F-F^{\prime})\in\mathrm{H}^{1}_{0}(I_{N})\cap\mathcal{H}_{N}$; in particular, $P_{N}(F-F^{\prime})$ belongs to the form domain $\mathrm{H}^{1}_{\alpha}(I_{N};\Gamma)\cap\mathcal{H}_{N}$, and consequently,
$\int_{I_{N}}\overline{\nabla\Psi(x)}\bm{\cdot}\nabla(F-F^{\prime})(x)+v(\rho_{\Psi,P_{N}(F-F^{\prime})})+w(\rho_{2,\Psi,P_{N}(F-F^{\prime})})-\lambda\langle\Psi,F-F^{\prime}\rangle$
$=\int_{I_{N}}\frac{1}{N!}\sum_{\sigma\in\mathcal{P}_{N}}\mathrm{sgn}(\sigma)\overline{\nabla(\sigma_{\#}\Psi)(x)}\bm{\cdot}\nabla(F-F^{\prime})(x)\mathrm{d}x+v(\rho_{\Psi,P_{N}(F-F^{\prime})})+w(\rho_{2,\Psi,P_{N}(F-F^{\prime})})-\lambda\langle\Psi,P_{N}(F-F^{\prime})\rangle$
$=a_{v,w}(\Psi,P_{N}(F-F^{\prime}))-\lambda\langle\Psi,P_{N}(F-F^{\prime})\rangle=0,$
@@
$\eta_{\epsilon}(x)\coloneqq\eta(d_{V}(x)/\epsilon),\quad\mbox{where}\quad d_{V}(x)\coloneqq\inf_{y\in V}||x-y||,$
@@
We now show that the Neumann trace is meaningful along any open subset of the boundary where $\Psi$ vanishes. The key to prove this is the locality of the form associated to $H_{N}(v,w)$.
@@
Without loss of generality, let us assume that $U\subset\subset E_{0}\coloneqq\{x=(x_{1},x^{\prime})\in\partial I_{N}:x_{1}=0,\quad x^{\prime}\in I_{N-1}\}$. Then, we let $\beta\in C^{\infty}_{c}(\mathbb{R};\mathbb{R})$ be a standard mollifier, i.e., $0\leq\beta\leq 1$, $\mathrm{supp}(\beta)\subset[-1,1]$ and $\beta=1$ on a neighborhood of 00, and define the function $F$ as
@@
$0=a_{v,w}(\Psi,P_{N}F_{\epsilon})-\lambda\langle\Psi,P_{N}F_{\epsilon}\rangle=$ $\int_{I_{N}}\left(1-\eta_{\epsilon}\right)\overline{\nabla\Psi}\bm{\cdot}\nabla F\mathrm{d}x+v(\rho_{\Psi,P_{N}F_{\epsilon}})+w(\rho_{2,\Psi,P_{N}F_{\epsilon}})-\lambda\langle\Psi,F_{\epsilon}\rangle$ (6.4)
$-\int_{I_{N}}F(x)\overline{\nabla\Psi(x)}\bm{\cdot}\nabla\eta_{\epsilon}(x)\mathrm{d}x,$ (6.5)
@@
$\lim_{\epsilon\downarrow 0^{+}}\rho_{\Psi,P_{N}F_{\epsilon}}=\rho_{\Psi,P_{N}F}\quad\mbox{in $\mathrm{H}^{1}(I)$}\quad\mbox{and}\quad\lim_{\epsilon\downarrow 0^{+}}\rho_{2,\Psi,P_{N}F_{\epsilon}}=\rho_{2,\Psi,P_{N}F}\quad\mbox{in $\mathrm{W}^{-1,p}(I_{2})$.}$
@@
$\partial_{\nu}\Psi(f)=\lim_{\epsilon\downarrow 0^{+}}\int_{I_{N}}F(x)\overline{\nabla\Psi(x)}\bm{\cdot}\nabla\eta_{\epsilon}(x)\mathrm{d}x,$
@@
$\partial_{\nu}\Psi(f)$ $=\lim_{\epsilon\downarrow 0^{+}}\int_{I_{N}}F(x)\overline{\nabla\Psi(x)}\bm{\cdot}\nabla\eta_{\epsilon}(x)\mathrm{d}x$
@@
Next, we recall that, by the definition of $\mathrm{H}^{1}_{\alpha}(I_{N})$ and antisymmetry, $\Psi$ satisfies
@@
Let $Jf$ and $Jf^{\prime}$ be extensions of $f$ and $f^{\prime}$ in $\mathrm{H}^{1}(I_{N})$. Then, recalling the proof of Lemma 5.7, we see that $P_{N}(Jf+Jf^{\prime})\in\mathrm{H}^{1}_{\alpha}(I_{N};\Gamma)\cap\mathcal{H}_{N}$. Therefore, by the definition of the Neumann trace, we have
$\partial_{\nu}\Psi(f)+\partial_{\nu}\Psi(f^{\prime})$ $=\partial_{\nu}\Psi(f+f^{\prime})=a_{v,w}\left(\Psi,P_{N}(Jf+Jf^{\prime})\right)-\lambda\langle\Psi,P_{N}(Jf+Jf^{\prime})\rangle=0.$
@@
Proof of Theorem 6.1 in the case of non-local BCs.
@@
$\partial_{\nu}\widetilde{\Psi}(f)$ $\coloneqq\widetilde{a}_{v,w}(\widetilde{\Psi},Jf)-\lambda\langle\widetilde{\Psi},Jf\rangle=a_{v,w}(T\widetilde{\Psi},TJf)-\lambda\langle T\widetilde{\Psi},TJf\rangle=\sqrt{N!}\partial_{\nu}\Psi(f)=0,$
@@
$\widetilde{a}_{v,w}(\widetilde{\Psi},\Phi)-\lambda\langle\widetilde{\Psi},\Phi\rangle=0,\quad\mbox{for any $\Phi\in Q_{\alpha}(W)$,}$ (6.9)
@@
$Q_{\alpha}(W)\coloneqq\{\Phi\in\mathrm{H}^{1}(S_{N};\Gamma_{N}\cup\Gamma_{\rm int}):(\gamma\Phi)(0,x^{\prime})-(-1)^{N-1}\alpha(\gamma\Phi)(x^{\prime},1)\quad\mbox{for $(0,x^{\prime})\in F_{0}\setminus W$}\},$
for any neighborhood $W\subset U\subset F_{0}$. Indeed, if $\Phi\in Q_{\alpha}(W)$, then $\Phi$ can be written as the sum of a function in the form domain of the reduced operator $\widetilde{H}^{\alpha}_{N}(v,w;\Gamma)$ and a function with trace supported in $\overline{W}\subset U$. To be precise, for any $\Phi\in Q_{\alpha}(W)$, we can define
@@
and note that $f_{\Phi}\in\mathrm{H}^{1/2}(\partial S_{N})$ because $f_{\Phi}$ is locally equal to $\gamma\Phi$ or a rigid motion of $\gamma\Phi$. Moreover, by the definition of $Q_{\alpha}(W)$, we have $\mathrm{supp}(\gamma\Phi-f_{\Phi})\subset\overline{W}$. Thus, since $Jf_{\Phi}$ belongs to the form domain of $\widetilde{H}_{N}^{\alpha}(v,w;\Gamma)$ (namely (5.13) intersected with $\mathrm{H}^{1}(S_{N};\Gamma_{N}\cap\partial S_{N})$), we have the desired decomposition $\Phi=(\Phi-Jf_{\Phi})+Jf_{\Phi}$, which proves the claim in (6.9).
Next, notice that (6.9) is equivalent to saying that $\widetilde{\Psi}$ is an eigenfunction of the operator associated to the form $\widetilde{a}_{v,w}$ but with the larger form domain $Q_{\alpha}(W)$. The important observation now is that the trace of functions in $Q_{\alpha}(W)$ have no restriction inside $W$. Therefore, we can repeat the extension argument in the proof of the case of local BCs. More precisely, we extend $S_{N}$ to $\Omega$ by adding an open ball whose intersection with $\partial S_{N}$ is compactly contained in $W$. Then $\widetilde{\Psi}$ extended by zero is also an eigenfunction of the extended operator $H_{\Omega}$ associated to (6.1) but this time with form domain
$Q_{\Omega}\coloneqq\{\Psi\in\mathrm{H}^{1}(\Omega;\Gamma):\Psi\rvert_{U\cap\partial\Omega}=0\quad\mbox{and}\quad(\gamma\Phi)(0,x^{\prime})-(-1)^{N-1}\alpha(\gamma\Phi)(x^{\prime},1)\quad\mbox{for $(0,x^{\prime})\in F_{0}\setminus W$}\}.$
Thus, recalling assumption (2.4), one can repeat the arguments in the proof of Lemma 5.7 to show that $Q_{\Omega}$ is positivity and reality preserving. In particular, Theorem 5.3 applies to $H_{\Omega}$ and therefore its ground-state $\Psi_{\Omega}$ is non-degenerate and almost everywhere strictly positive in $\Omega$. This in turn implies that $\Psi_{\Omega}$ is either orthogonal or parallel to $\widetilde{\Psi}$. However, neither of these options are possible because both functions are strictly positive in $S_{N}$ but $\widetilde{\Psi}$ vanishes identically in $\Omega\setminus S_{N}$ while $\Psi_{\Omega}$ does not. This yields the desired contradiction and completes the proof. ∎

### v2

@@ 5.1. Unique continuation along the boundary
Let $\Gamma\subset\partial I_{N}$, and let $\Psi$ be the unique ground-state of the self-adjoint realization of $\mathfrak{a}_{v,w}$ with form domain $\mathcal{Q}_{N}(\Gamma)=\mathrm{H}^{1}(I_{N};\Gamma)\cap\mathcal{H}_{N}$. Then $\gamma\Psi$ can not vanish identically on a relatively open set
@@
Moreover, if $\alpha\in\mathbb{R}\setminus\{0\}$ satisfies (2.5), then the same holds for the ground-state of the self-adjoint realization with form domain $\mathcal{Q}_{N}^{\alpha}(\Gamma)=\mathrm{H}^{1}_{\alpha}(I_{N};\Gamma)\cap\mathcal{H}_{N}$.
Proof of Theorem 5.1: case of local BCs.
@@
is continuous. We can thus define the extended self-adjoint operator $\widetilde{H}_{\Omega}$ via the sesquilinear form
$\widetilde{\mathfrak{a}}_{v,w,\Omega}(\Psi,\Phi)$ $\coloneqq\int_{\Omega\setminus S_{N}}\overline{\nabla\Psi(x)}\boldsymbol{\cdot}\nabla\Phi(x)\mathrm{d}x+\widetilde{\mathfrak{a}}_{v,w}(R\Psi,R\Phi)$
$=\int_{\Omega}\overline{\nabla\Psi(x)}\boldsymbol{\cdot}\nabla\Phi(x)\mathrm{d}x+(T^{\#}v)(R\Psi,R\Phi)+(T^{\#}w)(R\Psi,R\Phi),\quad\Psi,\Phi\in\mathrm{H}^{1}(\Omega;\Gamma_{S}).$ (5.1)
@@
$\widetilde{\mathfrak{a}}_{v,w,\Omega}(\widetilde{\Psi},\Phi)=\int_{\Omega}\overline{\nabla\widetilde{\Psi}}\boldsymbol{\cdot}\nabla\Phi\mathrm{d}x+(T^{\#}(v+w))(R\widetilde{\Psi},R\Phi)=\widetilde{\mathfrak{a}}_{v,w}(\widetilde{\Psi},R\Phi)=\lambda\langle\widetilde{\Psi},R\Phi\rangle=\lambda\langle\widetilde{\Psi},\Phi\rangle,$
where $\lambda$ is the ground-state energy of $\widetilde{H}_{N}(v,w;\Gamma)$. In particular, $\widetilde{\Psi}$ is an eigenfunction of $\widetilde{H}_{\Omega}$.
Next, notice that the pullback $R^{\#}V$ of any bounded multiplicative function $V\in\mathrm{L}^{\infty}(S_{N})$ is simply the extension of $V$ by zero to $\Omega$, so in particular, a bounded function in $\mathrm{L}^{\infty}(\Omega)$. Consequently, the pullback via $R$ of any sequence of bounded functions in $S_{N}$ approximating the quadratic form $T^{\#}(v+w)$ in $\mathrm{H}^{1}$ (as in (4.10)) yields a sequence of bounded functions approximating $R^{\#}T^{\#}(v+w)$. Consequently, the operator $\widetilde{H}_{\Omega}$ defined via (5.1) satisfies the assumptions of Theorem 4.6, and therefore, has a unique and a.e. strictly positive ground-state $\Psi_{\Omega}$.
This now yields a contradiction for the following reason. As we have shown that $\widetilde{\Psi}$ is an eigenfunction of $\widetilde{H}_{\Omega}$, it must be either orthogonal or proportional to $\Psi_{\Omega}$. However, $\widetilde{\Psi}$ cannot be orthogonal to $\Psi_{\Omega}$ as they are both non-negative (up to a global phase), and $\widetilde{\Psi}$ cannot be parallel to $\Psi_{\Omega}$ as the former vanishes on $\Omega\setminus S_{N}$ while the later is (a.e.) strictly positive in $\Omega$. This completes the proof. ∎
@@
Let $\alpha\in\mathbb{R}\setminus\{0\}$ and $\Psi\in\mathcal{Q}_{N}^{\alpha}(\Gamma)$ be an eigenfunction of $H_{N}^{\alpha}(v,w;\Gamma)$ with eigenvalue $\lambda$. Then for any $f\in\mathrm{H}^{1/2}(\partial I_{N})$, we define the Neumann trace $\partial_{\nu}\Psi$ as
$(\partial_{\nu}\Psi)(f)=\int_{I_{N}}\overline{\nabla\Psi(x)}\boldsymbol{\cdot}\nabla F(x)+v(\rho_{\Psi,P_{N}F})+w(\rho^{(2)}_{\Psi,P_{N}F})-\lambda\langle\Psi,F\rangle$ (5.2)
@@
$(P_{N}F)(x)=\frac{1}{N!}\sum_{\sigma\in\mathcal{P}_{N}}\mathrm{sgn}(\sigma)F(\sigma x)$ (5.3)
It is not hard to see that $\partial_{\nu}\Psi(f)$ is independent of the extension of $f$. Indeed, if we have two $\mathrm{H}^{1}$-extensions $F$ and $F^{\prime}$, then $F-F^{\prime}\in\mathrm{H}^{1}_{0}(I_{N})$, and therefore $P_{N}(F-F^{\prime})\in\mathrm{H}^{1}_{0}(I_{N})\cap\mathcal{H}_{N}$; in particular, $P_{N}(F-F^{\prime})$ belongs to the form domain $\mathcal{Q}_{N}^{\alpha}(\Gamma)$, and consequently,
$\int_{I_{N}}\overline{\nabla\Psi(x)}\boldsymbol{\cdot}\nabla(F-F^{\prime})(x)+v(\rho_{\Psi,P_{N}(F-F^{\prime})})+w(\rho^{(2)}_{\Psi,P_{N}(F-F^{\prime})})-\lambda\langle\Psi,F-F^{\prime}\rangle$
$=\int_{I_{N}}\frac{1}{N!}\sum_{\sigma\in\mathcal{P}_{N}}\mathrm{sgn}(\sigma)\overline{\nabla(\sigma_{\#}\Psi)(x)}\boldsymbol{\cdot}\nabla(F-F^{\prime})(x)\mathrm{d}x+v(\rho_{\Psi,P_{N}(F-F^{\prime})})+w(\rho^{(2)}_{\Psi,P_{N}(F-F^{\prime})})-\lambda\langle\Psi,P_{N}(F-F^{\prime})\rangle$
$=\mathfrak{a}_{v,w}(\Psi,P_{N}(F-F^{\prime}))-\lambda\langle\Psi,P_{N}(F-F^{\prime})\rangle=0,$
@@
[…]
@@
We now show that the Neumann trace is meaningful along any open subset of the boundary where $\Psi$ vanishes. The key to prove this is the locality of the form $\mathfrak{a}_{v,w}$.
@@
Without loss of generality, let us assume that $U\subset\subset E_{0}\coloneqq\{x=(x_{1},x^{\prime})\in\partial I_{N}:x_{1}=0,\quad x^{\prime}\in I_{N-1}\}$. Then, we let $\beta\in C^{\infty}_{c}(\mathbb{R};\mathbb{R})$ be a standard mollifier, i.e., $0\leq\beta\leq 1$, $\mathrm{supp}(\beta)\subset[-1,1]$ and $\beta=1$ on a neighborhood of 0, and define the function $F$ as
@@
$0=\mathfrak{a}_{v,w}(\Psi,P_{N}F_{\epsilon})-\lambda\langle\Psi,P_{N}F_{\epsilon}\rangle=$ $\int_{I_{N}}\left(1-\eta_{\epsilon}\right)\overline{\nabla\Psi}\boldsymbol{\cdot}\nabla F\mathrm{d}x+v(\rho_{\Psi,P_{N}F_{\epsilon}})+w(\rho^{(2)}_{\Psi,P_{N}F_{\epsilon}})-\lambda\langle\Psi,F_{\epsilon}\rangle$ (5.5)
$-\int_{I_{N}}F(x)\overline{\nabla\Psi(x)}\boldsymbol{\cdot}\nabla\eta_{\epsilon}(x)\mathrm{d}x,$ (5.6)
@@
$\lim_{\epsilon\downarrow 0^{+}}\rho_{\Psi,P_{N}F_{\epsilon}}=\rho_{\Psi,P_{N}F}\quad\mbox{in $\mathrm{H}^{1}(I)$}\quad\mbox{and}\quad\lim_{\epsilon\downarrow 0^{+}}\rho^{(2)}_{\Psi,P_{N}F_{\epsilon}}=\rho^{(2)}_{\Psi,P_{N}F}\quad\mbox{in $\mathrm{W}^{-1,p}(I_{2})$.}$
@@
$\partial_{\nu}\Psi(f)=\lim_{\epsilon\downarrow 0^{+}}\int_{I_{N}}F(x)\overline{\nabla\Psi(x)}\boldsymbol{\cdot}\nabla\eta_{\epsilon}(x)\mathrm{d}x,$
@@
$\partial_{\nu}\Psi(f)$ $=\lim_{\epsilon\downarrow 0^{+}}\int_{I_{N}}F(x)\overline{\nabla\Psi(x)}\boldsymbol{\cdot}\nabla\eta_{\epsilon}(x)\mathrm{d}x$
@@
Next, we recall that, by the definition of $\mathrm{H}^{1}_{\alpha}(I_{N})$ and antisymmetry, any $\Psi\in\mathcal{Q}_{N}^{\alpha}$ satisfies
@@
Let $Jf$ and $Jf^{\prime}$ be extensions of $f$ and $f^{\prime}$ in $\mathrm{H}^{1}(I_{N})$. Then, recalling the proof of Lemma 4.10, we see that $P_{N}(Jf+Jf^{\prime})\in\mathcal{Q}_{N}^{\alpha}(\Gamma)$. Therefore, by the definition of the Neumann trace, we have
$\partial_{\nu}\Psi(f)+\partial_{\nu}\Psi(f^{\prime})$ $=\partial_{\nu}\Psi(f+f^{\prime})=\mathfrak{a}_{v,w}\left(\Psi,P_{N}(Jf+Jf^{\prime})\right)-\lambda\langle\Psi,P_{N}(Jf+Jf^{\prime})\rangle=0.$
@@
Proof of Theorem 5.1: the case of non-local BCs.
@@
$\partial_{\nu}\widetilde{\Psi}(f)$ $\coloneqq\widetilde{\mathfrak{a}}_{v,w}(\widetilde{\Psi},Jf)-\lambda\langle\widetilde{\Psi},Jf\rangle=\mathfrak{a}_{v,w}(T\widetilde{\Psi},TJf)-\lambda\langle T\widetilde{\Psi},TJf\rangle=\sqrt{N!}\partial_{\nu}\Psi(f)=0,$
@@
$\widetilde{\mathfrak{a}}_{v,w}(\widetilde{\Psi},\Phi)-\lambda\langle\widetilde{\Psi},\Phi\rangle=0,\quad\mbox{for any $\Phi\in\mathcal{Q}_{W}$,}$ (5.10)
@@
$\mathcal{Q}_{W}\coloneqq\{\Phi\in\mathcal{Q}_{N}(\Gamma;S_{N}):(\gamma\Phi)(0,x^{\prime})-(-1)^{N-1}\alpha(\gamma\Phi)(x^{\prime},1)\quad\mbox{for $(0,x^{\prime})\in F_{0}\setminus W$}\},$
for any neighborhood $W\subset U\subset F_{0}$. Indeed, if $\Phi\in\mathcal{Q}_{W}$, then $\Phi$ can be written as the sum of a function in the form domain of the reduced operator $\widetilde{H}^{\alpha}_{N}(v,w;\Gamma)$ and a function with trace supported in $\overline{W}\subset U$. To be precise, for any $\Phi\in\mathcal{Q}_{W}$, we can define
@@
and note that $f_{\Phi}\in\mathrm{H}^{1/2}(\partial S_{N})$ because $f_{\Phi}$ is locally equal to $\gamma\Phi$ or a rigid motion of $\gamma\Phi$. Moreover, by the definition of $\mathcal{Q}_{W}$, we have $\mathrm{supp}(\gamma\Phi-f_{\Phi})\subset\overline{W}$. Thus, since $Jf_{\Phi}$ belongs to the form domain of $\widetilde{H}_{N}^{\alpha}(v,w;\Gamma)$ (namely (4.13) intersected with $\mathrm{H}^{1}(S_{N};\Gamma_{N}\cap\partial S_{N})$), we have the desired decomposition $\Phi=(\Phi-Jf_{\Phi})+Jf_{\Phi}$, which proves the claim in (5.10).
Next, notice that (5.10) is equivalent to saying that $\widetilde{\Psi}$ is an eigenfunction of the operator associated to the form $\widetilde{\mathfrak{a}}_{v,w}$ but with the larger form domain $\mathcal{Q}_{W}$. The important observation now is that the trace of functions in $\mathcal{Q}_{W}$ have no restriction inside $W$. Therefore, we can repeat the extension argument in the proof of the case of local BCs. More precisely, we extend $S_{N}$ to $\Omega$ by adding an open ball whose intersection with $\partial S_{N}$ is compactly contained in $W$. Then $\widetilde{\Psi}$ extended by zero is also an eigenfunction of the extended operator $\widetilde{H}_{\Omega}$ associated to (5.1) but this time with form domain
$\mathcal{Q}_{\Omega}\coloneqq\{\Psi\in\mathrm{H}^{1}(\Omega;\Gamma):\Psi\rvert_{U\cap\partial\Omega}=0\quad\mbox{and}\quad(\gamma\Phi)(0,x^{\prime})-(-1)^{N-1}\alpha(\gamma\Phi)(x^{\prime},1)\quad\mbox{for $(0,x^{\prime})\in F_{0}\setminus W$}\}.$
Thus, recalling assumption (2.5), one can repeat the arguments in the proof of Lemma 4.10 to show that $\mathcal{Q}_{\Omega}$ is positivity and reality preserving. In particular, Theorem 4.6 applies to $\widetilde{H}_{\Omega}$ and therefore its ground-state $\Psi_{\Omega}$ is non-degenerate and almost everywhere strictly positive in $\Omega$. This in turn implies that $\Psi_{\Omega}$ is either orthogonal or parallel to $\widetilde{\Psi}$, which gives us a contradiction as these functions can not be orthogonal since they are strictly positive in $S_{N}$ but can not be parallel because $\widetilde{\Psi}$ vanishes identically in $\Omega\setminus S_{N}$ while $\Psi_{\Omega}$ does not. ∎